OpenStudy (anonymous):

The perimeter of a rectangle is 50 cm and the length is 4 cm longer than the width. Write and solve the system of equations using the substitution method. A. width = 10 cm, length= 14 cm B. width = 10 1/2 cm, length= 14 1/2 cm C. width = 11 cm, length= 15 cm D. width = 23 cm, length= 27 cm

OpenStudy (anonymous):

D

OpenStudy (anonymous):

can u help with 2 ore

OpenStudy (anonymous):

YA

OpenStudy (anonymous):

Use a system of equations to solve the word problem. The sum of two numbers is 24. Their difference is 14. Find the numbers. A. 5 & 19 B. 7 & 17 C. 10 & 14 D. 6 & 20

OpenStudy (anonymous):

d

OpenStudy (anonymous):

can i get an opinion from another to make sure

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Solve the system, using substitution. 3a - b = 11 and 2a + 3b = 0 A. (3, 2) B. (3, -2) C. (-3, 2) D. (-3, -2) E. no solution

OpenStudy (campbell_st):

Perimeter is P length = l and width = w P = 2l + 2w so 50 = 2l + 2w then you know after dividing by 2 25 = l + w this is 1 equation and you know length is 4 longer than width l - w = 4 this is the other equation so there are your 2 equations. now you can solve the equations

OpenStudy (anonymous):

it got B. (3, -2)

OpenStudy (campbell_st):

I forgot to mention, D isn't the answer

OpenStudy (anonymous):

i got b

OpenStudy (campbell_st):

for the 1st question...D isn't the answer

OpenStudy (anonymous):

the what is it the other one said it was i just dont want to get it wrong/

OpenStudy (campbell_st):

well if you choose D to the 1st question you posted.... you did get it wrong

OpenStudy (anonymous):

i did not say d the other on in this chat said it was d i don't WANT TO GET OT WRONG ?

OpenStudy (campbell_st):

so go and read my post it has a method for finding the correct solution....

OpenStudy (anonymous):

do u knwo the answer please i have to go in 2 minutes

OpenStudy (campbell_st):

Open study isn't about providing answers... its about helping understanding...

OpenStudy (anonymous):

just never mind then yall is going to get me wrong and i dont have the time 4 that !!